Lasix 100 mg by gastrostomy tube once a day. Available: Lasix 40 mg per 5 ml.
Answer #1

ANSWER.

* Calculation of volume(ml) of drug required.

       Formula

            Volume(ml) = Desired dose(D)/Dose in hand(H) * quantity(Q).

   Here

       Desired dose(D) = 100 mg.

       Dose in hand(H) = 40 mg.

            Quantity(Q) = 5 ml.

As per above formula

     Volume(ml) of drug required = D/H * Q

                                                    = 100/40 * 5

                                                    = 2.5 * 5

                                                    = 12.5 ml.